OpenStudy (anonymous):

integral 1 / ( x^3 * sqrt (x^2 - 1))

OpenStudy (watchmath):

\(u=\sqrt{x^2-1}\). Then the integral is equal to \(\int\frac{du}{(u^2+1)^2}\,du\). Then subsitute \(u=\tan t\) to solve that :).